# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ Appendix C Proof of Theorem 2
$=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\mathbb{P}_{\mathcal{T}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (58)
$\geq\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Big{[}\alpha({\mathbf{x}})\Big{]}=1-\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}\Bigg{]},$ (59)
where eq. 59 follows from Theorem 1.121212To be explicit, in the series of equations above, we can go from eq. 58 to eq. 59 because $\mathcal{X}_{L}$ has zero measure. For brevity, let $t:=\frac{n\mu^{2}}{4}(1-2p)^{2}$. Also, we will omit the subscript ${\mathbf{x}}\sim\mathcal{D}$ henceforth for conciseness. Now, for any $\nu\in(0,1)$:
$\mathbb{E}\Big{[}\exp\Big{(}-t\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}\Big{]}=\underbrace{\mathbb{E}\Big{[}\exp\Big{(}-t\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}\Big{|}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{]}}_{\leq 1}\mathbb{P}\Big{(}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}\\ +\underbrace{\mathbb{E}\Big{[}\exp\Big{(}-t\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}\Big{|}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}>\nu\Big{]}}_{\leq\exp(-t\nu^{2})}\underbrace{\mathbb{P}\Big{(}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}>\nu\Big{)}}_{\leq 1}.$ (60)
Thus,
$\mathbb{E}\Big{[}\exp\Big{(}-t\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}\Big{]}\leq\mathbb{P}\Big{(}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}+\exp(-t\nu^{2}).$ (61)
From Lemma 3, we have for $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$:
$\mathbb{P}\Big{(}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}$ $\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}.$ (62)
Using this in eq. 61 and plugging in the value of $t$, we get:
$\mathbb{E}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}\Bigg{]}$ $\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\nu^{2}\Bigg{)},$ (63)
for any $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$. Since $\gamma<\frac{1}{2}$, we can choose $\nu=\frac{\gamma}{\sqrt{3d}}$. Plugging this choice into eq. 63 yields:
$\mathbb{E}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}\Bigg{]}$ $\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{12d}\Bigg{)}.$ (64)
Plugging this into eq. 59 gives us:
@@
$\text{acc}(\widehat{\bm{\theta}}_{0})$ $\geq 1-\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}-\exp\Bigg{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{12d}\Bigg{)}$ (65)
$=1-\exp\big{(}-\Theta(d)\big{)}-\exp\Bigg{(}-\frac{n{\mu^{2}}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Bigg{)}.$ (66)
(b) Upper bound. We have:
$\text{acc}(\widehat{\bm{\theta}}_{0})$ $=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Bigg{[}\mathbbm{1}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (67)
$=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\mathbb{P}_{\mathcal{T}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (68)
$\leq\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Big{[}\widehat{\alpha}({\mathbf{x}})\Big{]}=1-\frac{1}{3\sqrt{2\pi}}\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}\Bigg{]},$ (69)
where eq. 69 follows from Theorem 1.131313To be explicit, in the series of equations above, we can go from eq. 68 to eq. 69 because $\mathcal{X}_{L}$ has zero measure. For brevity, let $t:=(1+\sqrt{n}\mu(1-2p))^{2}$. Also, we will omit the subscript ${\mathbf{x}}\sim\mathcal{D}$ henceforth for conciseness. Now, for any $\nu>0$:
$\mathbb{E}\Big{[}\exp\big{(}-t\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}\big{)}\Big{]}$ $\geq\mathbb{E}\Big{[}\exp\big{(}-t\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}\big{)}\Big{|}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{]}\mathbb{P}\Big{(}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}$ (70)
$\geq\exp\big{(}-{t\nu^{2}}\big{)}\mathbb{P}\Big{(}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}.$ (71)
Let us choose $\nu=\frac{2\sqrt{2}\mu}{\sqrt{d-1}}$. Then, using the result of Lemma 4, we have:
$\mathbb{E}\Big{[}\exp\big{(}-t\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}\big{)}\Big{]}$ $\geq\frac{1}{2}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}\exp\Big{(}-\frac{8\mu^{2}t}{d-1}\Big{)}.$ (72)
Using this in eq. 69 after plugging in the value of $t=(1+\sqrt{n}\mu(1-2p))^{2}\leq 2(1+n\mu^{2}(1-2p)^{2})$, we get:
$\text{acc}(\widehat{\bm{\theta}}_{0})$ $\leq 1-\frac{1}{6\sqrt{2\pi}}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}\exp\Bigg{(}-\frac{16\mu^{2}(1+n\mu^{2}(1-2p)^{2})}{d-1}\Bigg{)}$ (73)
$\leq 1-\Theta(1)\exp\Bigg{(}-\frac{n\mu^{4}(1-2p)^{2}}{\Theta(d)}\Bigg{)}.$ (74)
This completes the proof. ∎
Lemma 3.
In the setting of the proof of Theorem 2, we have for any $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\leq\nu\Big{)}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}.$
Proof.
We have:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\geq\nu\Big{)}$ $=\mathbb{P}\Bigg{(}\frac{\big{(}{\mathrm{x}}^{(1)}\big{)}^{2}}{\big{(}{\mathrm{x}}^{(1)}\big{)}^{2}+\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}\big{)}^{2}}\geq\nu^{2}\Bigg{)}$ (75)
$=\mathbb{P}\Bigg{(}\sum_{j=2}^{d}\Big{(}\frac{{\mathrm{x}}^{(j)}}{\sigma}\Big{)}^{2}\leq\Big{(}\frac{1-\nu^{2}}{\nu^{2}}\Big{)}\Big{(}\frac{{\mathrm{x}}^{(1)}}{\sigma}\Big{)}^{2}\Bigg{)}.$ (76)
Let ${\mathrm{z}}:=\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}/\sigma\big{)}^{2}$ and $\overline{\mathrm{x}}^{(1)}:={\mathrm{x}}^{(1)}/\sigma$. Note that ${\mathrm{z}}$ is a chi-squared random variable with $(d-1)$ degrees of freedom and $\big{|}\overline{\mathrm{x}}^{(1)}\big{|}\geq\gamma$ per our setting. Thus, continuing from eq. 76, we have:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\geq\nu\Big{)}$ $\geq\mathbb{P}\Bigg{(}{\mathrm{z}}<\Big{(}\frac{1-\nu^{2}}{\nu^{2}}\Big{)}\gamma^{2}\Bigg{)}.$ (77)
A simple application of Theorem 1 of Ghosh, (2021) yields:
$\mathbb{P}\big{(}{\mathrm{z}}\geq 3(d-1)\big{)}\leq\exp\Bigg{(}-\frac{(d-1)}{2}(2-\log 3)\Bigg{)}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}.$ (78)
Thus, for any $w\geq 3(d-1)$, $\mathbb{P}({\mathrm{z}}<w)\geq 1-\exp\Big{(}-\frac{9(d-1)}{20}\Big{)}$. Using this in eq. 77 yields:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\geq\nu\Big{)}\geq 1-\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)},$ (79)
when $\Big{(}\frac{1-\nu^{2}}{\nu^{2}}\Big{)}\gamma^{2}\geq 3(d-1)$ or $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$. So:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\leq\nu\Big{)}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)},$ (80)
for $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$. This finishes the proof. ∎
Lemma 4.
In the setting of the proof of Theorem 2, we have for any $\nu\geq\frac{2\sqrt{2}\mu}{\sqrt{d-1}}$:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cot}(\beta({\mathbf{x}}))|\leq\nu\Big{)}\geq\frac{1}{2}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}.$
Proof.
We have:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cot}(\beta({\mathbf{x}}))|\leq\nu\Big{)}$ $=\mathbb{P}\Bigg{(}\frac{\big{(}{\mathrm{x}}^{(1)}\big{)}^{2}}{\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}\big{)}^{2}}\leq\nu^{2}\Bigg{)}$ (81)
$=\mathbb{P}\Bigg{(}\Big{(}\frac{{\mathrm{x}}^{(1)}}{\sigma}\Big{)}^{2}\leq\nu^{2}\sum_{j=2}^{d}\Big{(}\frac{{\mathrm{x}}^{(j)}}{\sigma}\Big{)}^{2}\Bigg{)}.$ (82)
Let ${\mathrm{z}}:=\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}/\sigma\big{)}^{2}$ and $\overline{\mathrm{x}}^{(1)}:={\mathrm{x}}^{(1)}/\sigma$. Note that ${\mathrm{z}}$ is a chi-squared random variable with $(d-1)$ degrees of freedom. Thus, continuing from eq. 82, we have:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cot}(\beta({\mathbf{x}}))|\leq\nu\Big{)}$ $\geq\mathbb{P}\Big{(}\big{(}\overline{\mathrm{x}}^{(1)}\big{)}^{2}\leq\nu^{2}{\mathrm{z}}\Big{)}$ (83)
$\geq\mathbb{P}\Bigg{(}\big{(}\overline{\mathrm{x}}^{(1)}\big{)}^{2}\leq\nu^{2}{\mathrm{z}}\text{ }\big{|}\text{ }{\mathrm{z}}>\frac{d-1}{2}\Bigg{)}\mathbb{P}\Bigg{(}{\mathrm{z}}>\frac{d-1}{2}\Bigg{)}$ (84)
$\geq\mathbb{P}\Bigg{(}\big{(}\overline{\mathrm{x}}^{(1)}\big{)}^{2}\leq\nu^{2}\Big{(}\frac{d-1}{2}\Big{)}\Bigg{)}\mathbb{P}\Bigg{(}{\mathrm{z}}>\frac{d-1}{2}\Bigg{)}.$ (85)
From Theorem 2 of Ghosh, (2021), we have that:
$\mathbb{P}\Bigg{(}{\mathrm{z}}>\frac{d-1}{2}\Bigg{)}\geq 1-\exp\Bigg{(}-\frac{d-1}{16}\Bigg{)}.$ (86)
Plugging this into eq. 85, we get:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cot}(\beta({\mathbf{x}}))|\leq\nu\Big{)}$ $\geq\mathbb{P}\Bigg{(}\big{(}\overline{\mathrm{x}}^{(1)}\big{)}^{2}\leq\nu^{2}\Big{(}\frac{d-1}{2}\Big{)}\Bigg{)}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}$ (87)
$=\mathbb{P}\Bigg{(}\big{|}\overline{\mathrm{x}}^{(1)}\big{|}\leq\nu\sqrt{\frac{d-1}{2}}\Bigg{)}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}.$ (88)
As per our setting, we have that $\mathbb{E}[\big{|}\overline{\mathrm{x}}^{(1)}\big{|}]=\mu$. Note that for $\nu\geq\frac{2\sqrt{2}\mu}{\sqrt{d-1}}$, we have that:
$\mathbb{P}\Bigg{(}\big{|}\overline{\mathrm{x}}^{(1)}\big{|}\leq\nu\sqrt{\frac{d-1}{2}}\Bigg{)}\geq\mathbb{P}\Big{(}\big{|}\overline{\mathrm{x}}^{(1)}\big{|}\leq 2\mu\Big{)}\geq\frac{1}{2},$ (89)
where the last step above follows from Markov’s inequality. Plugging in eq. 89 into eq. 88 gives us the desired result. ∎
Appendix D Proof of Theorem 3
Proof.
[…]
We first follow a standard argument to “reduce” the classification problem to a multi-way hypothesis testing problem. Let $\rho\in(0,1)$ be an arbitrary but fixed value which we can choose. We define a $\rho$-packing of $S^{d-1}$ as a set $\Theta=\{\bm{\theta}_{1},\dotsc,\bm{\theta}_{M}\}\subset S^{d-1}$ such that $\langle\bm{\theta}_{l},\bm{\theta}_{k}\rangle\leq\rho$ for $l\neq k$. Also define the $\rho$-packing number of $S^{d-1}$ as
$M(\rho,S^{d-1}):=\sup\{M\in\mathbb{N}:\text{there exists a $\rho$-packing $\Theta$ of $S^{d-1}$ with size $M$}\}\,.$ (90)
For convenience, we define the misclassification error of a classifier $\widehat{\bm{\theta}}\in S^{d-1}$ as $\textup{err}(\widehat{\bm{\theta}};\bm{\theta}^{*})=1-\textup{acc}(\widehat{\bm{\theta}};\bm{\theta}^{*})$. By our assumption,
$\delta\geq\sup_{\bm{\theta}^{\ast}\in S^{d-1}}\textup{err}(\widehat{\bm{\theta}};\bm{\theta}^{*})\geq\sup_{\bm{\theta}^{\ast}\in\Theta}\textup{err}(\widehat{\bm{\theta}};\bm{\theta}^{*})\,.$ (91)
In order to further lower bound the right hand side, we let $I$ be a random variable uniformly distributed on the hypothesis set $\{1,2,\dotsc,M\}$ and consider the case of $\bm{\theta}^{\ast}=\bm{\theta}_{I}$. We also define $\widehat{I}$ as the index of the element in $\Theta$ with maximum inner product with $\widehat{\bm{\theta}}$ (it does not matter how we break ties). We then have
$\sup_{\bm{\theta}^{\ast}\in\Theta}\textup{err}(\widehat{\bm{\theta}};\bm{\theta}^{*})$ $\geq\max_{i\in[M]}\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)$

### v2

@@ B.2 Lower bound $\widetilde{\alpha}_{0}(\mathbf{x})$
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)=y)\geq 1-\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{8\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,,$
which completes the derivation of $\widetilde{\alpha}_{0}(\bm{x})$.
Appendix C Proof of Theorem 4.2
Using Theorem 4.1 we have:
${\sf acc}({\widehat{\bm{\theta}}}_{0})$ $\leq\operatorname{\mathbb{E}}[\alpha_{0}(\bm{x})]$
$=1-\frac{1}{2\sqrt{2\pi}}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]$
$\leq 1-\frac{1}{2\sqrt{2\pi}}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\bigg{|}\frac{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\leq v^{2}\right]\mathbb{P}\left(\frac{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\leq v^{2}\right)$
$\leq 1-\frac{1}{2\sqrt{2\pi}}\exp\left(-5(1+\sqrt{n}(1-2p))^{2}v^{2}\right)\mathbb{P}\left(\frac{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\leq v^{2}\right)\,,$ (C.1)
for any value of $v>0$. We next lower bound the probability on the right-hand side. We have
$\mathbb{P}\left(\frac{|\langle\bm{x},{\bm{{\mu}}}\rangle|}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\leq v\right)$ $=\mathbb{P}\left(\frac{|y(1+u)\gamma^{2}|}{\left\|\bm{\Sigma}{\bm{z}}\right\|_{\ell_{2}}}\leq v\right)$
$\geq\mathbb{P}\left(\frac{(1+u)\gamma^{2}}{\lambda_{\min}\left\|\mathcal{P}^{\perp}_{\bm{\mu}}{\bm{z}}\right\|_{\ell_{2}}}\leq v\right)$
$=\mathbb{P}\left(\frac{(1+u)\gamma^{2}}{\left\|\mathcal{P}^{\perp}_{\bm{\mu}}{\bm{z}}\right\|_{\ell_{2}}}\leq v\lambda_{\min}\right)\,,$ (C.2)
where the first step holds since under our data model we have $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}$ and $\bm{\Sigma}^{1/2}\bm{x}=y(1+u)\bm{\Sigma}^{1/2}{\bm{{\mu}}}+\bm{\Sigma}\bm{z}=\bm{\Sigma}\bm{z}$, given that $\bm{\Sigma}^{1/2}{\bm{{\mu}}}=0$. Here $\mathcal{P}^{\perp}_{\bm{\mu}}$ is the projection onto the orthogonal space of ${\bm{{\mu}}}$.
Continuing from (C.2) we write
$\mathbb{P}\left(\frac{(1+u)\gamma^{2}}{\left\|\mathcal{P}^{\perp}_{\bm{\mu}}{\bm{z}}\right\|_{\ell_{2}}}\leq v\lambda_{\min}\right)$ $\geq\mathbb{P}\left((1+u)\gamma^{2}\leq v\lambda_{\min}\sqrt{d/2},\;\left\|\mathcal{P}^{\perp}_{\bm{\mu}}{\bm{z}}\right\|_{\ell_{2}}\geq\sqrt{d/2}\right)$
$=\mathbb{P}\left((1+u)\gamma^{2}\leq v\lambda_{\min}\sqrt{d/2}\right)\mathbb{P}\left(\left\|\mathcal{P}^{\perp}_{\bm{\mu}}{\bm{z}}\right\|_{\ell_{2}}\geq\sqrt{d/2}\right)\,,$ (C.3)
given that ${\bm{z}}$ and $u$ are independent. Next note that $\mathcal{P}^{\perp}_{\bm{\mu}}{\bm{z}}$ is distributed as a Gaussian vector in a $(d-1)$-dimensional space, and therefore $\left\|\mathcal{P}^{\perp}_{\bm{\mu}}\bm{z}\right\|_{\ell_{2}}^{2}$ is $\chi^{2}$ distribution with $(d-1)$-degree of freedom. We next use the following result about the tail bound of $\chi^{2}$ distribution to control $\left\|\mathcal{P}^{\perp}_{\bm{\mu}}\bm{z}\right\|_{\ell_{2}}$.
Lemma C.1
(Ghosh,, 2021, Theorem 4) Suppose $X\sim\chi^{2}_{d}(\lambda)$. Then for $0<\eta<d+\lambda$, we have
$\mathbb{P}(X<d+\lambda-\eta)\leq\exp\left[-\frac{d\eta^{2}}{4(d+2\lambda)^{2}}\right]\,.$
As a corollary of Lemma C.1, by choosing $\eta=\lambda+d/2$, we obtain $\mathbb{P}(X<d/2)\leq e^{-d/16}$ for any $\lambda\geq 0$. Using this we have
$\mathbb{P}\left(\left\|\mathcal{P}^{\perp}_{\bm{\mu}}\bm{z}\right\|_{\ell_{2}}^{2}<d/2\right)\leq e^{-d/16}\,.$ (C.4)
We next lower bound the other term on the right-hand side of (C.3). By Markov’s inequality for any non-negative random variable $X$, we have $\mathbb{P}(X\leq 2\operatorname{\mathbb{E}}[X])\geq 1/2$. Also $\operatorname{\mathbb{E}}[(1+u)\gamma^{2}]\leq 2\gamma^{2}$, since $\operatorname{\mathbb{E}}[u]\leq 1$ and therefore, by choosing $v=\frac{4\sqrt{2}\gamma^{2}}{\lambda_{\min}\sqrt{d}}$, we get
$\mathbb{P}\left((1+u)\gamma^{2}\leq v\lambda_{\min}\sqrt{d/2}\right)\geq 1/2\,.$ (C.5)
Combining (C.4), (C.5), (C.3) and (C.2) with the bound (C.1), we arrive at
@@
${\sf acc}({\widehat{\bm{\theta}}}_{0})\leq 1-\frac{1}{4\sqrt{2\pi}}\exp\left(-160(1+\sqrt{n}(1-2p))^{2}\frac{\gamma^{4}}{\lambda_{\min}^{2}d}\right)\left(1-e^{-d/16}\right)\,.$
To derive the lower bound on ${\sf acc}({\widehat{\bm{\theta}}}_{0})$ note that by using Theorem 4.1 we have
${\sf acc}({\widehat{\bm{\theta}}}_{0})$ $\geq\operatorname{\mathbb{E}}[\widetilde{\alpha}_{0}(\bm{x})]$